Listed below are the lead concentrations in μ​g/g measured in different traditional medicines. Use a 0.01 significance level to test the claim that the mean lead concentration for all such medicines is less than 16 μ​g/g.

18.5 17   10   4.5  15   16 11   19.5 11.5   15.5

The test statistic is ___________

The p value is ____________
Answer #1

Ans:

sample mean=13.85

sample standard deviation=4.571

n=10

Test statistic:

t=(13.85-16)/(4.571/SQRT(10))

t=-1.487

df=10-1=9

p-value=P(t<-1.487)=tdist(1.487,9,1)=0.0855

As,p-value>0.01,we fail to reject the null hypothesis.There is not sufficient evidence to support that claim that the mean lead concentration for all such medicines is less than 16 μ​g/g.